# Developing Instruction-Following Speech Language Model Without Speech Instruction-Tuning Data

### v1

@@ I Introduction
Developing SLMs presents significant challenges, particularly in constructing suitable speech-text training pairs to bridge the gap between text and speech modalities. While training strategies vary across studies, most require a substantial speech instruction-tuning stage to enable broad understanding of speech tasks. One common approach is utilizing textual metadata or labels from existing speech datasets, often employing human annotators to create task-specific instructions [7, 12] or using text-based LLMs to generate open-ended question-answer pairs [6, 5, 8, 9, 11, 16]. However, these approaches have severe limitations. The quality of the resulting dataset heavily depends on the annotators or system prompts, inevitably introducing annotation bias or ambiguity into the training data, such as specific output formats or word patterns. Consequently, the model must not only learn to understand speech features but also contend with annotation biases. We believe this could potentially lead to task overfitting [7] in SLMs, where the model becomes overly specialized in producing specific forms of output or performing particular tasks. Furthermore, given that text-based LLMs already possess the ability to understand and follow diverse instructions, it’s worth questioning whether a massive speech instruction-tuning stage is truly necessary for developing SLMs. We should consider if it’s possible to enhance the model’s speech comprehension capabilities without compromising the original versatility of the LLM. In this work, we demonstrate that the answer is ”yes”. Our approach addresses two critical challenges simultaneously: mitigating the forgetting problem often associated with specialized finetuning and reducing the need for extensive human effort in data preparation.
In this work, we introduce a dataset construction approach that combines two innovations: First, we leverage the text-based LLM in our SLM to generate speech-text pairs using textual metadata collected from speech. This crucial design minimizes the textual mismatch between the original LLM and new speech-text training data, allowing the model to focus primarily on learning speech information. These speech metadata covers a wide range of speech attributes, which are either from human annotation or extracted by several specialized speech models. Second, motivated by the idea of learning general speech concepts from speech captions [17], we employ a single prompt—”What can you hear from the audio?” for both data construction and model training, which significantly reduces the need for task-specific annotation. We call this carefully designed descriptive speech-text alignment approach DeSTA2.
@@ III-A Dataset construction
As depicted in Figure 1(Left), we encapsulate all relevant information from the speech into a structured format termed a ”seed transcript”. This format integrates text transcriptions with paralinguistic attributes, following a pattern such as: ”[00:00:00-00:00:03] How are you? (Gender: Female, Emotion: Happy…)”. Subsequently, we input the seed transcript into the LLM alongside the prompt ”What can you hear from the audio?” to generate comprehensive speech descriptions. These generated responses to the input context then serve as our training target for the end-to-end SLM model.
@@ V-D Examples on complex reasoning ability
Q: Analyze the audio, then answer ”What is the emotion of the speaker?”
Ours: Based on the audio, the speaker’s emotion is ”happy”. The
@@
Ours: Let’s break it down step by step:
@@ 1. The speech lasts for 2 seconds: [00:00:00 - 00:00:02]
2. The number of words is:”Thank you for calling. Goodbye.”(5 words)

### v2

# DeSTA2: Developing Instruction-Following Speech Language Model Without Speech Instruction-Tuning Data

@@ I Introduction
Developing SLMs presents significant challenges, particularly in constructing suitable speech-text training pairs to bridge the gap between text and speech modalities. While training strategies vary across studies, most require a substantial speech instruction-tuning stage to enable broad understanding of speech tasks. One common approach is utilizing textual metadata or labels from existing speech datasets, often employing human annotators to create task-specific instructions [7, 12] or using text-based LLMs to generate open-ended question-answer pairs [6, 5, 8, 9, 11, 16]. However, these approaches have severe limitations. The quality of the resulting dataset heavily depends on the annotators or system prompts, inevitably introducing annotation bias or ambiguity into the training data, such as specific output formats or word patterns. Consequently, the model must not only learn to understand speech features but also contend with annotation biases. We believe this could potentially lead to task overfitting [7] in SLMs, where the model becomes overly specialized in producing specific forms of output or performing particular tasks. Furthermore, given that text-based LLMs already possess the ability to understand and follow diverse instructions, it’s worth questioning whether a massive speech instruction-tuning stage is truly necessary for developing SLMs. We should consider if it’s possible to enhance the model’s speech comprehension capabilities without compromising the original versatility of the LLM. In this work, we demonstrate that the answer is "yes". Our approach addresses two critical challenges simultaneously: mitigating the forgetting problem often associated with specialized finetuning and reducing the need for extensive human effort in data preparation.
In this work, we introduce a dataset construction approach that combines two innovations: First, we leverage the text-based LLM in our SLM to generate speech-text pairs using textual metadata collected from speech. This crucial design minimizes the textual mismatch between the original LLM and new speech-text training data, allowing the model to focus primarily on learning speech information. These speech metadata covers a wide range of speech attributes, which are either from human annotation or extracted by several specialized speech models. Second, motivated by the idea of learning general speech concepts from speech captions [17], we employ a single prompt—"What can you hear from the audio?" for both data construction and model training, which significantly reduces the need for task-specific annotation. We call this carefully designed descriptive speech-text alignment approach DeSTA2.
@@ III-A Dataset construction
As depicted in Figure 1(Left), we encapsulate all relevant information from the speech into a structured format termed a "seed transcript". This format integrates text transcriptions with paralinguistic attributes, following a pattern such as: "[00:00:00-00:00:03] How are you? (Gender: Female, Emotion: Happy…)". Subsequently, we input the seed transcript into the LLM alongside the prompt "What can you hear from the audio?" to generate comprehensive speech descriptions. These generated responses to the input context then serve as our training target for the end-to-end SLM model.
@@ V-D Examples on complex reasoning ability
Q: Analyze the audio, then answer "What is the emotion of the speaker?"
DeSTA2: Based on the audio, the speaker’s emotion is "happy". The
@@
DeSTA2: Let’s break it down step by step:
@@ 1. The speech lasts for 2 seconds: [00:00:00 - 00:00:02]
2. The number of words is:"Thank you for calling. Goodbye."(5 words)
@@ VI Conclusion
VII Acknowledgment
The authors thank NVIDIA Taiwan AI R&D Center for the TRDC budget support and contributions to this research.